# Statistical Inference with Limited Memory: A Survey

### v1

@@ 6.4 Additional Results
The finite-memory simple hypothesis testing problem had an impact on other problems in the memory constrained literature. In [103], Hellman and Cover solved the two armed bandit with time-invariant finite memory problem by using a similar construction and proof elements from [84]. Horos and Hellman [104] allowed a confidence to be associated with each decision, with errors being weighted according to the confidence with which they are made. They found the $\varepsilon$-optimal class of rules to be deterministic. Flower and Hellman [105] examined the finite sample problem for Bernoulli case. They found that randomization was needed on all transitions toward the center state, i.e., on transitions from states of low uncertainty to states with higher uncertainty. Continuing the exploration of the finite-sample hypothesis testing with finite memory problem, Cover et al. [72] established the existence of an optimal rule and found its structure for time-varying algorithms, and characterised the optimal error probability of the $2$-state finite-sample problem that distinguishes between two Gaussian distributions with a different mean.
@@ 7 Parameter Estimation
The parameter estimation with finite-memory problem under the mean squared error loss function was first addressed by Roberts and Tooley [106], who tackled the problem of estimation under quadratic risk for a random variable with additive noise using time-varying machines. They restricted their rules to be of a special form which, although not optimal in general, does make sense. This restriction is that larger observations cause transitions to higher numbered states. In some problems (for example, the Cauchy distribution), very large observations yield very little information, and such a rule seems to be distinctly suboptimal. However, for Gaussian statistics, large observations are very informative, in which case their form does make sense. Koplowitz and Roberts [107] demonstrated necessary and sufficient conditions for the optimal state transition function.
@@
Wagner [92] uses rules similar to Cover’s [70] to estimate the mean of a distribution. For Bernoulli observations, Wagner’s scheme is very close to optimal, since its maximum absolute error is at most $\varepsilon$, with $S=O(1/\varepsilon)$ memory states. He achieved this result by partitioning the observations into blocks $\{S_{k}^{j},R_{k}^{j}\}_{k=1}^{\infty}$ sequentially testing whether $p>p_{j}$ or $p<p_{j}$, for $1\leq j\leq r$, and $0=p_{0}<p_{1}<\ldots<p_{r}=1,r=O(1/\varepsilon)$. By keeping the index $j$ of the favourite hypothesis and choosing the block lengths to follow the Borrel-Cantelli convergence conditions, the correct block is identified with probability $1$. This is an example of the methodology mentioned in Section 5.2, of breaking down a complex problem into multiple simple binary decision problems; we will see this again in the sequel.
@@
Hellman [108] examined the problem of estimating the mean $\mu$ of a Gaussian Distribution $\mathcal{N}(\mu,\sigma^{2})$ under the mean squared error loss, where $\mu$ is drawn from a known prior distribution $p(\mu)$. To obtain a lower bound, consider the simpler quantization problem, where we pick the best $S$-level scalar quantizer $Q(\mu)$ for $\mu$, minimizing the mean squared error $\operatorname{\mathbb{E}}(Q(\mu)-\mu)^{2}$. Clearly, this mean squared error is a lower bound for the mean squared error attained by any $S$-state estimation algorithm. One might assume that this quantization lower bound is not tight, however, quite surprisingly, Hellman showed that the quantization lower bound can be approached arbitrarily closely by the following (deterministic) construction: Let $\mu_{1}^{*}<\mu_{2}^{*}<\ldots<\mu_{S}^{*}$ be the optimal quantization points of an $S$-level quantizer. Consider the machine which transits from state $i$ to state $i+1$ whenever $X_{n}-(\mu_{i}^{*}+\mu_{i+1}^{*})/2\geq M$, and from state $i$ to state $i-1$ when $X_{n}-(\mu_{i-1}^{*}+\mu_{i}^{*})/2\leq-M$, and which estimates $\mu$ to be $\mu_{i}^{*}$ when it is in state $i$. Thus, the ratio between transition probabilities toward states with better $\mu$ estimates and transition probabilities toward states with worse $\mu$ estimates behaves exponentially with $M$, so that, as $M\rightarrow\infty$, the stationary distribution approaches a singleton on $\mu^{*}$. Thus, in the Gaussian case, the mean estimation problem can be reduced to a quantization problem, that is, for any prior $p(\mu)$ on $\mu$ the minimum quadratic risk attained by an $S$-state estimator is equal to the quadratic risk attained by the optimal $S$-level quantizer.
@@
Samaniego [109] initiated the work on the problem of estimating the parameter of a coin using time-invariant finite-memory machines. He considered a Bayesian setup in which a prior is placed on the parameter distribution, and the loss function is a generalized quadratic loss function of the form $(\hat{\theta}-\theta)^{2}\theta^{\beta-1}(1-\theta)^{\gamma-1}$, for some $\beta,\gamma\geq 0$. Furthermore, he restricted to machines that can only make transitions between adjacent states, which he refers to as tridiagonal algorithms. Within the class of all tridiagonal algorithms, a particular rule was shown to be minimax as well as locally Bayes w.r.t the uniform prior for $S\leq 30$. Leighton and Rivest [110] followed up the work on estimating the parameter of a coin, and have found an exact characterization of the smallest attainable mean squared error for this problem. Specifically, they proved that
@@
Samaniego’s construction from [109] is analyzed and the mean squared error is recovered for any $\theta$ and $S$. Essentially, this finite-state algorithm uses $S$ states to estimate the number of ones in a sliding window of length $S-1$, thus it is commonly referred to in the literature as ”Imaginary Sliding Window” (See Figure 3). In this construction, if the machine is in state $i$ and the fresh sample is ”$1$”, we move to state $i+1$ with probability $1-\frac{i}{S-1}$ or stay in state $i$ with probability $\frac{i}{S-1}$. If the fresh sample is ”00”, we move to state $i-1$ with probability $\frac{i}{S-1}$ or stay in state $i$ with probability $1-\frac{i}{S-1}$. The estimate function is $\hat{\theta}(i)=\frac{i}{S-1}$.
@@
The key ingredient in the proof is the use of the Markov Chain Tree Theorem (see the elegant proof of Anantharam and Tsoucas [111]), which implies that the limiting distribution $\pi(j)$ of an $S$-state Ergodic Markov chain (which is a unique stationary distribution) can be expressed as ratios of $S-1$ degree polynomials with non-negative coefficients. [^4] The theorem actually holds in the general case, for which the limiting distribution might not exist, with $\pi_{j}$ replaced by $\bar{p}_{1j}$, where $\bar{p}_{1j}$ is the long-run average transition probability that the chain will be in state $j$ given that it started in state $1$. Specifically, it is the $(1,j)$ element of the matrix $\bar{P}=\lim_{n\rightarrow\infty}n^{-1}\sum_{i=1}^{n}P^{i-1}$, where $P$ is the transition matrix of the chain, and $\bar{P}$ always exists as $P$ is stochastic. Specifically, we have
@@ 7.1 Deterministic Algorithm for Bias Estimation
We consider now the bias estimation problem with restriction to deterministic machines. Following the results of Section 6, one might expect a degradation in performance when comparing to optimal randomized machines. In fact, in the same seminal work of [110], the authors further constructed a deterministic $S$-state estimation procedure by de-randomizing Samaniego’s construction, and as a result showed that $R_{\mathsf{det}}^{*}(S)=O(\log{S}/S)$. They then conjectured that this is indeed the deterministic minimax risk of the problem. A nice interpretation of their conjecture is the naturally appealing claim that an optimal deterministic algorithm can be obtained by derandomizing an optimal random algorithm. In their deterministic algorithm, randomness is extracted from the measurements via a Von Neumann like procedure [112], augmenting each state with $O(\log(S))$ additional states [^5] This derandomization is almost tight. We need to generate $S$ Bernoulli r.v.s with parameters $i/S$ for $i=1,\ldots,S$ for transition probabilities. Assume we use machines with $N_{t}$ states respectively. The blowup factor $N$ in the number of states is the average of $N_{t}$, and there must be at least $S/2$ machines with $N_{t}\leq 2N$. There are at most $(2N)^{4N+1}$ binary input FSM with at most $2N$ states, hence generating $S/2$ different Bernoulli parameters dictates that $(2N)^{4N+1}\geq S/2$, which implies $N=\Omega(\log(S)/\log\log(S))$.. After standing unresolved for a couple of decades, this conjecture was recently disproved by Berg et al. [87], who proved that
@@
which implies that the memory complexity of the problem is unaffected by the restriction to deterministic estimators.
@@
In particular, the state space is partitioned into $K$ disjoint sets denoted by $\mathcal{S}_{1},\ldots,\mathcal{S}_{K}$, where the estimation function value is the same inside each $\mathcal{S}_{i}$. These sets are referred to as mini-chains, where the $i$th mini-chain consists of $N_{i}$ states, and is designed to solve the composite binary hypothesis testing problem:
@@
Recall that in section 6.2 we saw that $O(K)$ states are needed to resolve between two hypothesis that are $1/K$ apart with constant error probability. Thus, letting $K=O(\sqrt{S})$, the described algorithm estimates $\theta$ to within $1/\sqrt{S}$ standard deviations using $S$ states, which implies the correct minimax risk.
@@ 7.2 Linear Regression
In [115], Sharan et al. considered the case of linear regression where the $X$ are drawn independently from a $d$-dimensional isotropic Gaussian and the noise is drawn uniformly from the interval $[-2^{-d/5},2^{-d/5}]$. They show that any algorithm with at most $d^{2}/4$ bits of memory requires at least $\Omega(d\log\log 1/\varepsilon)$ for $\varepsilon$ sufficiently small as a function of $d$. In contrast, for such $\varepsilon$, $\theta$ can be recovered to error $\varepsilon$ with probability $1-o(1)$ with memory $O(d^{2}\log(1/\varepsilon))$ bits using $d$ samples. The uniform small error is crucial for the proof, as it reframes the problem as a quantization problem that allows the authors to leverage branching programs for the lower bound.
@@ 7.3 Additional Results
In [120], Rajwan and Feder considered universal finite-memory machines for coding binary sequences, and gave lower bounds for deterministic, randomized, time-invariant or time-variant machines. Meron and Feder [121] showed that for universal prediction of worst case sequences under the square error loss, the optimal asymptotic expected redundancy is $O(1/S)$ and it is achieved by the construction of [110]. For the log-loss case, using the same construction results in an upper bound on the redundancy of $O(\log S/S)$, while Pinsker’s inequality gives a lower bound of $\Omega(1/S)$. In a follow up work [31], the same authors also consider universal prediction of worst case sequences using deterministic finite-state machines, proving a lower bound on the regret of $\Omega(S^{-4/5})$, and suggested the ”Exponentially Decaying Memory” machine, which achieves regret of $O(S^{-2/3})$.
@@ 8 Distribution Property Testing
In the uniformity testing problem we observe a sequence of independent identically distributed random variables drawn from an unknown distribution $P$ over $[k]$, which is guaranteed to be either the uniform distribution $U_{k}$, or $\varepsilon$-far from the uniform distribution under the total variation distance. In other words, there are two composite hypotheses in this problem, $\mathcal{F}_{0}=\{U_{k}\}$, and $\mathcal{F}_{1}$ which contains all the distributions $P$ that satisfy $d_{\mathsf{TV}}(P,U_{k})>\varepsilon$, and we want to identify the correct (composite) hypothesis with probability of success at least $2/3$. We note that this problem is a particular instance of the identity testing problem, in which $\mathcal{F}_{0}=\{Q\}$, for some known distribution $Q$ (not necessarily uniform). However, it was shown by Goldreich [122] that the problem of uniformity testing can be reduced to identity testing, and, in particular, that for every distribution $Q$ over $[k]$ and every $\varepsilon>0$, it holds that $\varepsilon$-testing identity to $Q$ reduces to $\varepsilon/3$-testing identity to $U_{6k}$, and that the same reduction can be used for all $\varepsilon>0$.
@@
The lower bound of [79] is proved in a Bayesian setting using information theoretic tools. Typically, for the purpose of lower bounds, the minimax setting is replaced by a Bayesian one, where the alternative hypothesis is the so-called Paninski prior or Paniniski mixture, a uniform mixture over all Paniniski distributions of the form (38) (the specific distribution from the mixture is picked only once, uniformly at random, and then we observe i.i.d. samples from it). A distribution from this mixture is constructed as follows: Let $X=2J-(Z_{J}\oplus B_{J})$, where $J\sim\mathsf{Unif}([k/2]),B\sim\mathsf{Bern}^{\otimes k/2}(1/2-\varepsilon)$, and $B$ is independent of $I$. Note that for $\varepsilon=0$, this results in the uniform distribution regardless of $Z$. On the other hand, for any choice of $z\in\{0,1\}^{k/2}$, each of these distributions is exactly $P_{\text{pan}}^{\varepsilon}(z)$, which satisfies the total variation condition $d_{\mathsf{TV}}(P_{\text{pan}}^{\varepsilon}(z),U_{k})=\varepsilon$, namely is a valid alternative to the uniform hypothesis.
The lower bound is established for the case where $Z$ is first drawn uniformly on $\{0,1\}^{k/2}$ and then $X_{1},\ldots,X_{n}$ are generated i.i.d. either from $P_{\text{pan}}^{\varepsilon}(Z)$ or from $U_{k}$. A high-level interpretation of the argument is as follows: For any algorithm, the mutual information between the memory content and $Z$ is at most $b$, so that, from standard rate-distortion results [9], the best estimate we can get for any $Z_{i}$ has error $h_{2}^{-1}(1-\frac{b}{k})=\frac{1}{2}-\Omega\left(\sqrt{\frac{b}{k}}\right)$ on average, where $h_{2}(\cdot)$ is the binary entropy function, and $h_{2}^{-1}$ is its inverse restricted to $[0,1/2]$. In each measurement we observe, the random variable $I$ entails no information on whether or not $B_{J}$ is biased, and hence the informative part of the observation is $Z_{J}\oplus B_{J}$, which is unbiased if $B_{J}\sim\mathsf{Bern}(1/2)$, corresponding to $P=U_{K}$, and has bias $\Omega\left(\varepsilon\cdot\sqrt{\frac{b}{k}}\right)$ on average, under the alternative hypothesis. Thus, assuming both hypotheses are equally likely a-priori, the mutual information for determining whether the bias of $B_{J}$ is 00 or $\varepsilon$, is $O(\varepsilon^{2}\cdot\frac{b}{k})$ per measurement, and $O(n\varepsilon^{2}\cdot\frac{b}{k})$ in total. Thus we need $n=\Omega(k/(b\varepsilon^{2}))$ samples to succeed.
@@
where the upper bound holds for $b>>\frac{\log k}{\varepsilon^{2}}$ bits. A natural question that arises from [79] is: to what extent can this memory size be reduced if we are allowed to process more samples? In a recent work, Berg et al. [88] considered the problem of finding the minimax memory complexity of uniformity testing, and proposed an algorithm that succeeds with $O(\log(k/\varepsilon))$ memory bits. Moreover, the authors gave a lower bound of $\Omega(\log(k))$ on the memory complexity. Next, we outline the derivation of these bounds.
@@
We now describe an algorithm that succeeds with $O\left(\frac{k\log k}{\varepsilon}\right)$ states. In a nutshell, the suggested algorithm reduces uniformity testing to a sequence of binary hypothesis tests. The underlying idea is that if $d_{\mathsf{TV}}(P,U_{k})>\varepsilon$, then there must be some symbol $i\in[k]$ such that either $\Pr(X=1|X\in\{1,i\})$ or $\Pr(X=i|X\in\{1,i\})$ are $\varepsilon$-biased, where by $\varepsilon$-bias we mean equals $1/2+\Omega(\varepsilon)$. On the other hand, if $P=U_{k}$, none of these probabilities is $\varepsilon$-biased, in fact, they are exactly $\mathsf{Bern}(1/2)$. Following this observation, we can construct a finite state machine whose state space is partitioned into $2k$ disjoint sets (or mini-chains) denoted by $\mathcal{S}_{i}$, which are identical binary hypothesis testing $\operatorname{\mathsf{RUNS}}(N,1/2+\Omega(\varepsilon),1/2)$ machines with $N=O(1/\varepsilon\cdot\log k)$ states. These mini-chains test, for all $i>1$, whether $\Pr(X=1|X\in\{1,i\})$ or $\Pr(X=i|X\in\{1,i\})$ are $\varepsilon$-biased. If some $\mathcal{S}_{i}$ tests positive for $\varepsilon$-bias, the algorithm decides that $P$ is $\varepsilon$-far from uniform and terminates, otherwise it moves to the next mini-chain, whereas if no test is positive for $\varepsilon$-bias, the machine decides that $P$ is uniform. Due to the properties of the $\operatorname{\mathsf{RUNS}}(N,p,q)$ algorithm, the fact that each machine has $N=O(1/\varepsilon\cdot\log k)$ states guarantees that Bernoulli parameters that have $O(\varepsilon)$ bias between them can be distinguished with probability of error $\ll 1/k$, thus all tests are successful with some constant error probability ($1/3$, for example). We note that the effective run time of the above algorithm is $O\left(k^{2}e^{O(1/\varepsilon)}\right)$. This follows since there are $O(k)$ mini-chains that we traverse in sequence, each is activated every $O(k)$ samples on average, and the mixing time of each one is $e^{O(1/\varepsilon)}$.

### v2

@@ 6.4 Additional Results
The finite-memory simple hypothesis testing problem had an impact on other problems in the memory constrained literature. In [108], Hellman and Cover solved the two armed bandit with time-invariant finite memory problem by using a similar construction and proof elements from [5]. Horos and Hellman [109] allowed a confidence to be associated with each decision, with errors being weighted according to the confidence with which they are made. They found the $\varepsilon$-optimal class of rules to be deterministic. Flower and Hellman [110] examined the finite sample problem for the Bernoulli case. They found that randomization was needed on all transitions toward the center state, i.e., on transitions from states of low uncertainty to states with higher uncertainty. Continuing the exploration of the finite-sample hypothesis testing with finite memory problem, Cover et al. [20] established the existence of an optimal rule, found its structure for time-varying algorithms, and characterised the optimal error probability of the $2$-state finite-sample problem that distinguishes between two Gaussian distributions with a different mean.
@@ 7 Parameter Estimation
The parameter estimation with finite-memory problem was first addressed by Roberts and Tooley [111], who tackled the problem of estimation under quadratic risk for a random variable with additive noise using time-varying machines. They worked under the restriction that larger observations cause transitions to higher numbered states. In some problems (for example, the Cauchy distribution), very large observations yield very little information, and such a rule seems to be distinctly suboptimal. However, for Gaussian statistics, large observations are very informative, in which case their form does make sense. Koplowitz and Roberts [112] later demonstrated necessary and sufficient conditions for the optimal state transition function.
@@
Wagner [97] used rules similar to Cover’s [4] to estimate the mean of a distribution. For Bernoulli observations, Wagner’s scheme is very close to optimal, since its maximum absolute error is at most $\varepsilon$, with $S=O(1/\varepsilon)$ memory states. He achieved this result by partitioning the observations into blocks $\{S_{k}^{j},R_{k}^{j}\}_{k=1}^{\infty}$ sequentially testing whether $p>p_{j}$ or $p<p_{j}$, for $1\leq j\leq r$, and $0=p_{0}<p_{1}<\ldots<p_{r}=1,r=O(1/\varepsilon)$. By keeping the index $j$ of the favourite hypothesis and choosing the block lengths to follow the Borrel-Cantelli convergence conditions, the correct block is identified with probability $1$. This is an example of the methodology mentioned in Section 5.2, of breaking down a complex problem into multiple simple binary decision problems; we will see this again in the sequel.
@@
Hellman [113] examined the problem of estimating the mean $\mu$ of a Gaussian Distribution $\mathcal{N}(\mu,\sigma^{2})$ under the mean squared error loss, where $\mu$ is drawn from a known prior distribution $p(\mu)$. To obtain a lower bound, consider the simpler quantization problem, where we pick the best $S$-level scalar quantizer $Q(\mu)$ for $\mu$, minimizing the mean squared error $\operatorname{\mathbb{E}}(Q(\mu)-\mu)^{2}$. Clearly, this provides a lower bound for the mean squared error attained by any $S$-state estimation algorithm. One might assume that this quantization lower bound is not tight, however, quite surprisingly, Hellman showed that the quantization lower bound can be approached arbitrarily closely by the following (deterministic) construction: Let $\mu_{1}^{*}<\mu_{2}^{*}<\ldots<\mu_{S}^{*}$ be the optimal quantization points of an $S$-level quantizer. Consider the machine which transits from state $i$ to state $i+1$ whenever $X_{n}-(\mu_{i}^{*}+\mu_{i+1}^{*})/2\geq M$, and from state $i$ to state $i-1$ when $X_{n}-(\mu_{i-1}^{*}+\mu_{i}^{*})/2\leq-M$, and which estimates $\mu$ to be $\mu_{i}^{*}$ when it is in state $i$. Thus, the ratio between transition probabilities toward states with better $\mu$ estimates and transition probabilities toward states with worse $\mu$ estimates behaves exponentially with $M$, so that, as $M\rightarrow\infty$, the stationary distribution approaches a singleton on $\mu^{*}$, the closest value to $\mu$. Thus, in the Gaussian case, the mean estimation problem can be reduced to a quantization problem, that is, for any prior $p(\mu)$ on $\mu$ the minimum quadratic risk attained by an $S$-state estimator is equal to the quadratic risk attained by the optimal $S$-level quantizer.
@@
Samaniego [114] initiated the work on the problem of estimating the parameter of a coin using time-invariant finite-memory machines. He considered a Bayesian setup in which a prior is placed on the parameter distribution, and the loss function is a generalized quadratic loss function of the form $(\hat{\theta}-\theta)^{2}\theta^{\beta-1}(1-\theta)^{\gamma-1}$, for some $\beta,\gamma\geq 0$. Furthermore, he restricted to machines that can only make transitions between adjacent states, which he refers to as tridiagonal algorithms. Within the class of all tridiagonal algorithms, a particular rule was shown to be minimax as well as locally Bayes w.r.t the uniform prior for $S\leq 30$. Leighton and Rivest [115] followed up the work on estimating the parameter of a coin, and found an exact characterization of the smallest attainable mean squared error for this problem. Specifically, they proved that
@@
Samaniego’s construction from [114] is analyzed and the mean squared error is recovered for any $\theta$ and $S$. Essentially, this finite-state algorithm uses $S$ states to estimate the number of ones in a sliding window of length $S-1$, thus it is commonly referred to in the literature as ”Imaginary Sliding Window” (See Figure 3). In this construction, if the machine is in state $i$ and the fresh sample is ”$1$”, we move to state $i+1$ with probability $1-\frac{i}{S-1}$ or stay in state $i$ with probability $\frac{i}{S-1}$. If the fresh sample is ”00”, we move to state $i-1$ with probability $\frac{i}{S-1}$ or stay in state $i$ with probability $1-\frac{i}{S-1}$. The estimate function is $\hat{\theta}(i)=\frac{i-1}{S-1}$ for $i\in[S]$.
@@
The key ingredient in the proof is the use of the Markov Chain Tree Theorem (see the elegant proof of Anantharam and Tsoucas [116]), which implies that the limiting distribution $\pi_{j}$ of an $S$-state ergodic Markov chain (which is a unique stationary distribution) can be expressed as ratios of $S-1$ degree polynomials with non-negative coefficients. [^5] The theorem actually holds in the general case, for which the limiting distribution might not exist, with $\pi_{j}$ replaced by $\bar{p}_{1j}$, where $\bar{p}_{1j}$ is the long-run average probability that the chain will be in state $j$ given that it started in state $1$. Specifically, it is the $(1,j)$ element of the matrix $\bar{P}=\lim_{n\rightarrow\infty}n^{-1}\sum_{i=1}^{n}P^{i-1}$, where $P$ is the transition matrix of the chain, and $\bar{P}$ always exists as $P$ is stochastic. Specifically, we have
@@ 7.1 Deterministic Algorithm for Bias Estimation
We consider now the bias estimation problem with restriction to deterministic machines. Following the results of Section 6, one might expect a degradation in performance when comparing to optimal randomized machines. In fact, in the same seminal work of [115], the authors further constructed a deterministic $S$-state estimation algorithm by de-randomizing Samaniego’s construction, and as a result showed that $R_{\mathsf{det}}^{*}(S)=O(\log{S}/S)$. They then conjectured that this is indeed the deterministic minimax risk of the problem. A nice interpretation of their conjecture is the naturally appealing claim that an optimal deterministic algorithm can be obtained by derandomizing an optimal random algorithm. In their deterministic algorithm, randomness is extracted from the measurements via a Von Neumann like procedure [117], augmenting each state with $O(\log(S))$ additional states [^6] This derandomization is almost tight. To simulate the algorithm, we need to generate $S$ Bernoulli r.v.s with parameters $i/S$ for $i=1,\ldots,S$ for transition probabilities. Assume we use machines with $N_{t}$ states respectively. The blowup factor $N$ in the number of states is the average of $N_{t}$, and there must be at least $S/2$ machines with $N_{t}\leq 2N$. There are at most $(2N)^{4N+1}$ binary input FSM with at most $2N$ states, hence generating $S/2$ different Bernoulli parameters dictates that $(2N)^{4N+1}\geq S/2$, which implies $N=\Omega(\log(S)/\log\log(S))$.. After standing unresolved for a couple of decades, this conjecture was recently disproved by Berg et al. [92], who proved that
@@
which implies that the memory complexity of the problem is unaffected by the restriction to deterministic algorithms.
@@
In particular, the state space is partitioned into $K=O(\sqrt{S})$ disjoint sets denoted by $\mathcal{S}_{1},\ldots,\mathcal{S}_{K}$, where the estimation function value is the same inside each $\mathcal{S}_{i}$. These sets are referred to as mini-chains, where the $i$th mini-chain consists of $N_{i}$ states, and is designed to solve the composite binary hypothesis testing problem:
@@
Recall that in section 6.2 we saw that $O(K)$ states are needed to resolve between two hypothesis that are $1/K$ apart with constant error probability. Thus, under the assignment $K=O(\sqrt{S})$, the described algorithm can estimate $\theta$ to within $1/\sqrt{S}$ standard deviations using $S$ states, which implies the correct minimax risk.
@@ 7.2 Linear Regression
In [120], Sharan et al. considered the case of linear regression where the $X$ are drawn independently from a $d$-dimensional isotropic Gaussian and the noise is drawn uniformly from the interval $[-2^{-d/5},2^{-d/5}]$. They show that any algorithm with at most $d^{2}/4$ bits of memory requires at least $\Omega(d\log\log 1/\varepsilon)$ samples for $\varepsilon$ sufficiently small as a function of $d$. In contrast, for such $\varepsilon$, $\theta$ can be recovered to error $\varepsilon$ with probability $1-o(1)$ with memory $O(d^{2}\log(1/\varepsilon))$ bits using $d$ samples. The uniform small error is crucial for the proof, as it reframes the problem as a quantization problem that allows the authors to leverage branching programs for the lower bound.
@@ 7.3 Additional Results
In [125], Rajwan and Feder considered universal finite-memory machines for coding binary sequences, and gave lower bounds for deterministic, randomized, time-invariant and time-varying machines. Meron and Feder [126] showed that for universal prediction of worst case sequences under the squared error loss, the optimal asymptotic expected redundancy is $O(1/S)$ and it is achieved by the construction of [115]. For the log-loss case, using the same construction results in an upper bound on the redundancy of $O(\log S/S)$, while Pinsker’s inequality gives a lower bound of $\Omega(1/S)$. In a follow up work [39], the same authors also consider universal prediction of worst case sequences using deterministic finite-state machines, proving a lower bound on the regret of $\Omega(S^{-4/5})$, and suggested the ”Exponentially Decaying Memory” machine, which achieves regret of $O(S^{-2/3})$.
@@ 8 Distribution Property Testing
In the uniformity testing problem we observe a sequence of independent identically distributed random variables drawn from an unknown distribution $P$ over $[k]$, which is guaranteed to be either the uniform distribution $U_{k}$, or $\varepsilon$-far from the uniform distribution under the total variation distance. In other words, there are two composite hypotheses in this problem, $\mathcal{F}_{0}=\{U_{k}\}$, and $\mathcal{F}_{1}$ which contains all the distributions $P$ that satisfy $d_{\mathsf{TV}}(P,U_{k})>\varepsilon$, and we want to identify the correct (composite) hypothesis with probability of success at least $2/3$. We note that this problem is a particular instance of the identity testing problem, in which $\mathcal{F}_{0}=\{Q\}$, for some known distribution $Q$ (not necessarily uniform). However, it was shown by Goldreich [127] that the problem of identity testing can be reduced to uniformity testing and, in particular, that for every distribution $Q$ over $[k]$ and every $\varepsilon>0$, it holds that $\varepsilon$-testing identity to $Q$ reduces to $\varepsilon/3$-testing identity to $U_{6k}$, and that the same reduction can be used for all $\varepsilon>0$.
@@
The lower bound of [85] is proved in a Bayesian setting using information theoretic tools. Typically, for the purpose of lower bounds, the minimax setting is replaced by a Bayesian one, where the alternative hypothesis is the so-called Paninski prior or Paniniski mixture, a uniform mixture over all Paniniski distributions of the form (38) (the specific distribution from the mixture is picked only once, uniformly at random, and then we observe i.i.d. samples from it). A distribution from this mixture is constructed as follows: Let $X=2J-(Z_{J}\oplus B_{J})$, where $J\sim\mathsf{Unif}([k/2]),B\sim\mathsf{Bern}^{\otimes k/2}(1/2-\varepsilon)$, and $B$ is independent of $J$. Note that for $\varepsilon=0$, this results in the uniform distribution regardless of $Z$. On the other hand, for any choice of $z\in\{0,1\}^{k/2}$, each of these distributions is exactly $P_{\text{pan}}^{\varepsilon}(z)$, which satisfies the total variation condition $d_{\mathsf{TV}}(P_{\text{pan}}^{\varepsilon}(z),U_{k})=\varepsilon$, namely is a valid alternative to the uniform hypothesis.
The lower bound is established for the case where $Z$ is first drawn uniformly on $\{0,1\}^{k/2}$ and then $X_{1},\ldots,X_{n}$ are generated i.i.d. either from $P_{\text{pan}}^{\varepsilon}(Z)$ or from $U_{k}$. A high-level interpretation of the argument is as follows: For any algorithm, the mutual information between the memory content and $Z$ is at most $b$, so that, from standard rate-distortion results [14], the best estimate we can get for any $Z_{i}$ has error $h_{2}^{-1}(1-\frac{b}{k})=\frac{1}{2}-\Omega\left(\sqrt{\frac{b}{k}}\right)$ on average, where $h_{2}(\cdot)$ is the binary entropy function, and $h_{2}^{-1}$ is its inverse restricted to $[0,1/2]$. In each measurement we observe, the random variable $J$ entails no information on whether or not $B_{J}$ is biased, and hence the informative part of the observation is $Z_{J}\oplus B_{J}$, which is unbiased if $B_{J}\sim\mathsf{Bern}(1/2)$, corresponding to $P=U_{K}$, and has bias $\Omega\left(\varepsilon\cdot\sqrt{\frac{b}{k}}\right)$ on average, under the alternative hypothesis. Thus, assuming both hypotheses are equally likely a-priori, the mutual information for determining whether the bias of $B_{J}$ is 00 or $\varepsilon$, is $O(\varepsilon^{2}\cdot\frac{b}{k})$ per measurement, and $O(n\varepsilon^{2}\cdot\frac{b}{k})$ in total. Thus we need $n=\Omega(k/(b\varepsilon^{2}))$ samples to succeed.
@@
where the upper bound holds for $b\gg\frac{\log k}{\varepsilon^{2}}$ bits. A natural question that arises from [85] is to what extent this memory size can be reduced if we are allowed to process more samples. Another open question posed by the authors is whether the problem can be solved with $o(\log k)$ memory bits. The first question was addressed by Meir [131], who showed that one can achieve the sample complexity of [85] whenever the memory size in bits is $\Omega(\log(k/\varepsilon))$. This is done via comparison graphs, which are a general model for batch collision estimators, whose memory need only store the samples used for collision testing and a global counter. The second open question was resolved in a recent work of Berg et al. [93] on finding the memory complexity of uniformity testing. In the paper, the authors gave a lower bound of $\Omega(\log k)$ on the memory complexity in bits, thus answering the question from [85] in the negative. They further proposed an algorithm that succeeds with $\log(k/\varepsilon)+o(\log\log k)$ memory bits. Next, we outline the derivation of these bounds.
@@
The authors describe an algorithm that succeeds with $O\left(\frac{k\log k}{\varepsilon}\right)$ states by reducing uniformity testing to a sequence of binary hypothesis tests (an approach that should be familiar to us by now). The underlying idea is that if $d_{\mathsf{TV}}(P,U_{k})>\varepsilon$, then there must be some symbol $i\in[k]$ such that either $\Pr(X=1|X\in\{1,i\})$ or $\Pr(X=i|X\in\{1,i\})$ are $\varepsilon$-biased, where by $\varepsilon$-bias we mean equals $1/2+\Omega(\varepsilon)$. On the other hand, if $P=U_{k}$, none of these probabilities is $\varepsilon$-biased, in fact, they are exactly $\mathsf{Bern}(1/2)$. Following this observation, we can construct a finite state machine whose state space is partitioned into $2k$ disjoint sets (or mini-chains) denoted by $\mathcal{S}_{i}$, which are identical binary hypothesis testing $\operatorname{\mathsf{RUNS}}(N,1/2+\Omega(\varepsilon),1/2)$ machines with $N=O(1/\varepsilon\cdot\log k)$ states. These mini-chains test, for all $i>1$, whether $\Pr(X=1|X\in\{1,i\})$ or $\Pr(X=i|X\in\{1,i\})$ are $\varepsilon$-biased. If some $\mathcal{S}_{i}$ tests positive for $\varepsilon$-bias, the algorithm decides that $P$ is $\varepsilon$-far from uniform and terminates, otherwise it moves to the next mini-chain, whereas if no test is positive for $\varepsilon$-bias, the machine decides that $P$ is uniform. Due to the properties of the $\operatorname{\mathsf{RUNS}}(N,p,q)$ algorithm, the fact that each machine has $N=O(1/\varepsilon\cdot\log k)$ states guarantees that Bernoulli parameters that have $O(\varepsilon)$ bias between them can be distinguished with probability of error $\ll 1/k$, thus all tests are successful with some constant error probability ($1/3$, for example). We note that the effective run time of the above algorithm is $O\left(k^{2}e^{O(1/\varepsilon)}\right)$. This follows since there are $O(k)$ mini-chains that we traverse in sequence, each is activated every $O(k)$ samples on average, and the mixing time of each one is $e^{O(1/\varepsilon)}$.